OpenStudy (anonymous):

Can somebody check this for me? Simplify. Do not use negative exponents in the answer. Z ^-6/ Z -^2 -Z*Z*Z*Z*Z*Z = Z^-6 -Z*Z= Z^-2 So Z^-6 / Z ^-2= Z^3

OpenStudy (anonymous):

you subtract exponents with like bases when dividing \[\frac{z^{-6}}{x^{-2}}=z^{-6-(-2)}=z^{-4}=\frac{1}{z^4}\]

OpenStudy (anonymous):

no, a negative exponent means reciprocal so x^-2 = 1/x^2

OpenStudy (anonymous):

thats supposed to be a z in the denominator of the 1st part

OpenStudy (anonymous):

it's actually 1/z^4 because when you divide exponents, you have to subtract them which would give you (-6)-(-2)which would be z^-4. when you have a z raised to a -exponent, you can write it as a fraction. therefore it's 1/z^4 :) hope that helps!!!

OpenStudy (anonymous):

\[x^{-2} = {1 \over x^2} \] \[x^{-100000} = {1 \over x^{100000}}\] and \[{1 \over x^{-2}} = x^2\]

OpenStudy (anonymous):

That helps tons! thank you all! Just needed to check my work but wasn't sure. I appreciate it!